# Atomistic modeling of the channeling process with account for ionizing collisions: Preliminary results

### v1

@@ Abstract
We present preliminary results of simulations of the channeling process with account for the ionizing collisions that accompany the passage of a charged projectile in a crystalline medium. The simulations have been performed by means of the MBN Explorer software package. taking into account that ionizing collisions are random, fast and local they can be incorporated into the classical molecular dynamics framework according to their probabilities. The description of this methodology is given in the paper. The case studies presented refer to channeling of 195, 270 & 855 MeV electrons and of 270 & 500 MeV positrons in single diamond crystals along the (111) (electrons) and (110) (positrons) planar directions.
1 Introduction
@@
In recent years, advanced atomistic computational modeling of the channeling process and the process of the photon emission as well as of other related phenomena beyond the continuous potential framework has been carried out using the multi-purpose computer package MBN Explorer [14, 15] and the special multi-task software toolkit with graphical user interface MBN Studio [16]. The computation accounts for the interaction of projectiles with separate atoms of the environments, whereas many different interatomic potentials implemented in MBN Explorer support rigorous simulations of various media [17, 18]. For simulations of the channeling and related phenomena, an additional module has been incorporated into MBN Explorer to compute the motion of relativistic projectiles along with dynamical simulations of the environments. This methodology, called relativistic MD, has been extensively applied to simulate passage of ultra-relativistic charged particles in oriented crystals accompanied by emission of intensive radiation. A comprehensive description of the multiscale all-atom relativistic molecular dynamics approach as well as a number of case studies related to modeling channeling and photon emission by ultra-relativistic projectiles are presented in a review article [19].
Additional feature that has been implemented in MBN Explorer and is discussed in Section 2 below concerns the impact of the ionizing collisions on the channeling properties of ultra-relativistic charged particles. The events of inelastic scattering of a projectile particle from individual atoms lead to a random change in the direction of the particle’s velocity resulting in an additional mechanism of the increase in its transverse energy. Rigorous treatment of the ionization process can only be achieved by means of quantum mechanics. However, taking into account that such events are random, fast and local they can be incorporated into the classical molecular dynamics framework according to their probabilities. Similar methodology has been already implemented in MBN Explorer within the framework of the irradiation-driven molecular dynamics [26]. Preliminary results obtained for several related case studies are presented in Section 3.
2 Methodology
The MBN Explorer package [14, 17] implements the method of relativistic classical molecular dynamics [15] to model the motion an ultra-relativistic particle in an external field or/and in an atomic environment. The code integrates the relativistic equations of motion of a particle of mass $m$, charge $q$ and energy $\varepsilon$ written in the following form (see, e.g., [20]):
@@ 2 Methodology
The second term in (4), ${\bf F}_{\rm rd}$, is a radiation damping force that accounts for the reaction of the emitted radiation on the motion of the charged particle, see, e.g. [20, 21]. Its action leads to a gradual decrease of the particle’s energy. For ultra-relativistic projectiles of very high energies (tens of GeV and above) this force becomes quite noticeable, so that it must be accounted for to properly simulate the motion, see recent experiments [22, 23, 24]. The methodology that has been developed to incorporate ${\bf F}_{\rm rd}$ into the relativistic molecular dynamics framework implemented in the MBN Explorer package is described in Ref.[25].
Applied to a charged particle’s passage through an atomic environment, the system (3) describes the classical motion of the particle in an electrostatic field of the medium’s atoms. In this case, both terms on the right-hand side of Eq. (5) are dependent on the field’s intensity which is is calculated as ${\bf E}=-\bm{\nabla}\phi({\bf r})$ with $\phi({{\bf r}})$ standing for the field’s potential equal to the sum of potentials of individual atoms
@@
where ${\bf r}_{j}$ denotes the position vector of the nucleus of the $j$th atom. For a neutral atom, the potential $\phi({\bf r})$ is a rapidly decreasing function at the distances larger than the average atomic radius. Therefore, at each step of integration of Eqs. (3), the sum in (6) can account only for the atoms inside the sphere of a specified cut-off radius $\rho$ with the center at ${\bf r}$. Typically, $\rho$ is chosen to be order of magnitude larger than the average atomic radius. The search for such atoms is carried out by means of the linked cell algorithm implemented in the MBN Explorer package. In the course of particle’s passage through a medium, the environment is generated dynamically so that the simulation box follows the moving particle. For a crystalline environment, the positions of the atoms are generated accounting for random displacement from the nodes due to thermal vibrations. More detailed description on the algorithms used to compute the trajectories at various scales, including macroscopically large ones, are presented in Refs. [14, 15, 17, 19, 25].
Classical equations of motion (3) do not account for random events of inelastic scattering of a projectile particle from individual atoms leading to the atom’s excitation or ionization. These events results in the two following changes in the particle motion:
  • Random decrease in the projectile energy (so-called, ionization losses).
  • Random change in the direction of the projectile motion characterized by two scattering angles $\theta$ and $\phi$ measured with respect to the instant velocity ${\bf v}$.
Rigorous treatment of the inelastic collision evens can only be achieved by means of quantum mechanics. However, taking into account that such events are random, fast and local they can be incorporated into the classical mechanics framework according to their probabilities. This approach has been already implemented in MBN Explorer within the framework of irradiation-driven molecular dynamics [26]. Similar methodology, described below, has been realized in the code to account for the ionizing collisions. Its implementation followed the scheme presented in Ref. [27].
For an ultra-relativistic particle passing through a medium, the probability ${\rm d}^{2}W$ of the relative energy loss ${\rm d}\Delta={\rm d}\varepsilon/\varepsilon$ over the path ${\rm d}s\approx c{\rm d}t$ due to ionizing collision with quasi-free electrons can be calculated using the following expression (see, e.g., Eqs. (1.7)-(1.7b) in [28] and §33.2.7 in [29]):
@@
Here $r_{0}=2.818\times 10^{-13}$ stands for the classical electron radius and $n_{\mathrm{e}}(\textbf{r})$ is the local volume density of electrons. In the points ${\bf r}\neq{\bf r}_{j}$, i.e. not at the nuclei, the density is related to the potential (6) via the Poisson equation:
@@
In a single ionizing collision, the value of relative energy loss $\Delta=(\varepsilon-\varepsilon^{\prime})/\varepsilon$, where $\varepsilon$ and $\varepsilon^{\prime}<\varepsilon$ are the particle’s energies before and after the collision, lies within the interval
@@
At each step $\Delta t$ of integration of Eqs. (3) an ionizing collision is treated as a probabilistic event. Once it occurs and the value $\Delta$ is determined, one calculates a round scattering angle $\theta$ measured with respect to the instant velocity ${\bf v}(t)$:
@@
where $L_{\rm ion}$ is a spatial path over which the probability of an ionizing collision with an arbitrary energy transfer is equal to one:
@@
Therefore, the probability of an ionizing collision with the energy transfer $\Delta$ to occur over the path $\Delta s=c\Delta t$ is given by
@@
To simulate the probability of an ionizing collision to take place, a uniform random deviate $x\in[0,1]$ is generated and compared with $\Delta s/L_{\rm ion}$. If $x\leq\Delta s/L_{\rm ion}$ then the collision happens and another random deviate $\Delta$ is generated and used in Eq. (14) to calculate $\theta$.
@@
To account for or to disregard the ionizing collisions the following schemes are implemented in MBN Explorer:
  • the local electron density $n_{\mathrm{e}}(\textbf{r})$ is calculated using Eq. (8) where atomic potentials $\phi_{\rm at}$ are considered within the frameworks of the Molière [30] or Pacios [31] approximations.
@@
  • to disregard the ionizing collisions one sets $n_{\mathrm{e}}(\textbf{r})\equiv 0$.
3 Case studies: Preliminary results
@@
The case studies presented below refer to the channeling of $\varepsilon=195,270,855$ MeV electrons and $\varepsilon=270,500$ MeV positrons in 200 and 400 microns thick single diamond crystals. For electrons, the aforementioned energy range corresponds that available at the Mainz Mikrotron (MAMI) facility. A number of channeling experiments with the electron beam have been carried out at MAMI with various crystal targets during the last one and a half decade [32, 33, 6, 34, 35, 36]. More recently, within the framework of the TECHNO-CLS project [13], a new 500 MeV positron beam line has been designed at MAMI [37] and the first experimental tests with the positron beam have been carried out [38, 39].
The aim of the present study was to quantify the impact of the ionizing collisions on the channeling efficiency of projectiles of different charge and energy passing through comparatively thick crystalline targets. For doing this we have followed general methodology implemented in MBN Explorer to generate particles’ trajectories in a crystalline environment that accounts for randomness in sampling the incoming projectiles (the key factors here are beam size, divergence and direction with respect to the crystal orientation) as well as in displacement of the lattice atoms from the nodal positions due to thermal vibrations (for more details see Refs. [15, 19]). As a result, each trajectory corresponds to a unique crystalline environment and, therefore, all simulated trajectories are statistically independent and can be analyzed further to quantify the channeling process.
For each case study approximately $N_{0}=1.2\times 10^{4}$ of trajectories have been simulated and analyzed. The initial analysis has been carried out in terms of the two channeling fractions: $f_{\rm ch,0}(z)=N_{\rm ch,0}(z)/N_{0}$ and $f_{\rm ch}(z)=N_{\rm ch}(z)/N_{0}$ [19]. Here $N_{\rm ch,0}(z)$ denotes for the number of particles that were accepted to the channeling mode at the crystal entrance ($z=0$) and move in the channel up to distance $z$ where they dechannel. The ratio ${\cal A}=N_{\rm ch,0}(0)/N_{0}$ defines the fraction of the incoming particles accepted at the entrance. The quantity $N_{\rm ch}(z)$ stands for the total number of particles which move in the channeling mode at the distance $z$.
At comparatively large distances $z$ from the entrance point the fractions $f_{\rm ch,0}(z)$ decreases following the exponential decay law $\exp(-z/L_{\rm d})$, where $L_{\rm d}$ is the dechanneling length [1, 40]. The fraction $f_{\rm ch}(z)$ can be approximated by means of the fitting formula [19]:

### v2

# Atomistic modeling of the channeling process with and without account for ionising collisions: A comparative study

@@ Abstract
This paper presents a quantitative analysis of the impact of inelastic collisions of ultra-relativistic electrons and positrons, passing through oriented crystalline targets, on the channeling efficiency and on the intensity of the channeling radiation. The analysis is based on the numerical simulations of the channeling process performed using the MBN Explorer software package. The ionising collisions, being random, fast and local events, are incorporated into the classical molecular dynamics framework according to their probabilities. This methodology is outlined in the paper. The case studies presented refer to electrons with energy $\varepsilon$ ranging from 270 to 1500 MeV and positrons with $\varepsilon=530$ MeV incident on thick (up to 1 mm) single diamond, silicon and germanium crystals oriented along the (110) and (111) planar directions. In order to elucidate the role of the ionising collisions, the simulations were performed with and without account for the ionising collisions. The case studies presented demonstrate that both approaches yield highly similar results for the electrons. For the positrons, the ionising collisions reduce significantly the channeling efficiency. However, it has been observed that this effect does not result in a corresponding change in the radiation intensity.
I Introduction
@@
In recent years, advanced atomistic computational modeling of the channeling process and the process of the photon emission as well as of other related phenomena beyond the continuous potential framework has been carried out using the multi-purpose computer package MBN Explorer [14, 15] and the special multi-task software toolkit with graphical user interface MBN Studio [16]. The computation accounts for the interaction of projectiles with separate atoms of the environments, whereas many different interatomic potentials implemented in MBN Explorer support rigorous simulations of various media [17, 18]. For simulations of the channeling and related phenomena, an additional module has been incorporated into MBN Explorer to compute the motion of relativistic projectiles along with dynamical simulations of the environments. This methodology, called relativistic molecular dynamics (MD), has been extensively applied to simulate passage of ultra-relativistic charged particles in oriented crystals accompanied by emission of intensive radiation. A comprehensive description of the multiscale all-atom relativistic MD approach as well as a number of case studies related to modeling channeling and photon emission by ultra-relativistic projectiles are presented in a review article [19].
The aim of the present study is to quantify the impact of the ionising collisions on the channeling efficiency and on the photon emission spectra for projectile particles of different charge and energy passing through comparatively thick crystalline targets. For doing this we have followed general methodology implemented in MBN Explorer to generate particles’ trajectories in a crystalline environment that accounts for randomness in sampling the incoming projectiles as well as in displacement of the lattice atoms from the nodal positions due to thermal vibrations. As a result, each trajectory corresponds to a unique crystalline environment. Another phenomenon that contributes to the statistical independence of the simulated trajectories, concerns the events of inelastic scattering of a projectile particle from the crystal atoms resulting in the excitation or ionisation of the atom. These collisions lead to a random change in the direction of the particle’s velocity, which in turn leads to a change in its transverse energy. These quantum events are random and occur on the atomic scale in terms of time and space, therefore they can be incorporated into the classical mechanics framework in accordance with their probabilities.
The methodology implemented in MBN Explorer to account for the ionising collisions is discussed in Section II. In Section III we present and discuss the results of simulations of the channeling process of the sub-GeV electrons and positrons in different crystalline targets. A summary of the results obtained is given in Section IV.
II Methodology
The MBN Explorer package [14, 17] implements the method of relativistic classical MD [15] to model the motion an ultra-relativistic particle in an external field or/and in an atomic environment. The code integrates the relativistic equations of motion of a particle of mass $m$, charge $q$ and energy $\varepsilon$ written in the following form (see, e.g., [20]):
@@ II Methodology
The second term in (4), ${\bf F}_{\rm rd}$, is a radiation damping force that accounts for the reaction of the emitted radiation on the motion of the charged particle, see, e.g. [20, 21]. Its action leads to a gradual decrease of the particle’s energy. For ultra-relativistic projectiles of very high energies (tens of GeV and above) this force becomes quite noticeable, so that it must be accounted for to properly simulate the motion, see recent experiments [22, 23, 24]. The methodology that has been developed to incorporate ${\bf F}_{\rm rd}$ into the relativistic MD framework implemented in the MBN Explorer package is described in Ref.[25].
Applied to the passage of a charged particle through an atomic environment, the environment, the system (3) describes the classical motion of the particle in an electrostatic field of the atoms of the medium. In this case, the two terms on the right-hand side of eq. (5) depend on the intensity of the field, which can be calculated as is calculated as ${\bf E}=-\bm{\nabla}\phi({\bf r})$ where $\phi({{\bf r}})$ is the potential of the field is equal to the sum of the potentials of the individual atoms.
@@
where ${\bf r}_{j}$ denotes the position vector of the nucleus of the $j$th atom. For a neutral atom, the potential $\phi({\bf r})$ is a rapidly decreasing function at distances greater than the average atomic radius. Therefore, at each integration step of eq. (3), the sum in (6) can only include the atoms inside the sphere of a given cut-off radius $\rho$ with the centre at ${\bf r}$. Typically $\rho$ is chosen to be an order of magnitude larger than the average atomic radius. The search for such atoms is done using the linked cell algorithm implemented in the MBN Explorer package. As the particle passes through a medium, the environment is generated dynamically, so that the simulation box follows the moving particle. For a crystalline environment the positions of the atoms are generated taking into account random displacement from the nodes due to thermal vibrations. More detailed description on the algorithms used to compute the trajectories at various scales, including macroscopically large ones, are presented in Refs. [14, 15, 17, 19, 25].
The classical equations of motion (3) do not take into account for events of inelastic scattering of a projectile particle from individual atoms, which lead to the excitation or ionisation of the atom. These events result in two subsequent changes in the particle motion:
  • A random decrease in the energy of the projectile (ionisation losses).
  • A random change in the direction of the projectile motion characterised by two scattering angles, $\theta$ and $\phi$, measured with respect to the instant velocity, ${\bf v}$.
Rigorous treatment of inelastic collision events can only be achieved by means of quantum mechanics. However, given the random, fast and local nature of such events, they can be incorporated into the classical mechanics framework according to their probabilities. This approach has been already implemented in MBN Explorer within the framework of irradiation-driven MD [26]. A similar methodology, described below, has been realised in the code to account for the ionising collisions. Its implementation followed the scheme presented in Ref. [27].
For an ultra-relativistic particle passing through a medium, the probability ${\rm d}^{2}W$ of the relative energy loss ${\rm d}\Delta={\rm d}\varepsilon/\varepsilon$ over the path ${\rm d}s\approx c{\rm d}t$ due to ionising collision with quasi-free electrons can be calculated using the following expression (see, e.g., Eqs. (1.7)-(1.7b) in [28] and §33.2.7 in [29]):
@@
Here $r_{0}=2.818\times 10^{-13}$ is the classical electron radius and $n_{\mathrm{e}}(\textbf{r})$ is the local volume density of electrons. In the points ${\bf r}\neq{\bf r}_{j}$, i.e. not at the nuclei, the density is related to the potential (6) via the Poisson equation:
@@
In a single collision, the value of relative energy loss $\Delta=(\varepsilon-\varepsilon^{\prime})/\varepsilon$, where $\varepsilon$ and $\varepsilon^{\prime}<\varepsilon$ are the particle’s energies before and after the collision, lies within the interval
@@
At each step $\Delta t$ of integration of Eqs. (3) an ionising collision is treated as a probabilistic event. Once it occurs and the value $\Delta$ is determined, one calculates a round scattering angle $\theta$ measured with respect to the instant velocity ${\bf v}(t)$:
@@
where $L_{\rm ion}$ is a spatial path over which the probability of an ionising collision with an arbitrary energy transfer is equal to one:
@@
Therefore, the probability of an ionising collision with the energy transfer $\Delta$ to occur over the path $\Delta s=c\Delta t$ is given by
@@
To simulate the probability of an ionising collision to take place, a uniform random deviate $x\in[0,1]$ is generated and compared with $\Delta s/L_{\rm ion}$. If $x\leq\Delta s/L_{\rm ion}$ then the collision happens and another random deviate $\Delta$ is generated and used in Eq. (14) to calculate $\theta$.
@@
To account for or to disregard the ionising collisions the following schemes are implemented in MBN Explorer:
  • the local electron density $n_{\mathrm{e}}(\textbf{r})$ is calculated using Eq. (8) where atomic potentials $\phi_{\rm at}$ are considered within the frameworks of the Molière [30] or Fernandes Pacios [31] approximations.
@@
  • to disregard the ionising collisions one sets $n_{\mathrm{e}}(\textbf{r})\equiv 0$.
III Case studies
@@
The case studies presented below refer to the channeling of $\varepsilon=270,855$ and 1500 MeV electrons (e-) and 530 MeV positrons (e+) in single diamond (C), silicon (Si) and germanium (Ge) crystals oriented, with respect to the incident beam, along the (110) and (111) planar directions. The beams were assumed to be ideally collimated. The values of crystal thickness $L$ along the beam used in the simulations are (i) 200, 200, 400 microns for 270, 855, 1500 MeV electrons, respectively, and (ii) 1000 microns for the positrons. For electrons, the aforementioned energy range corresponds that available at the Mainz Mikrotron (MAMI) facility [32]. A number of channeling experiments with the electron beam have been carried out at MAMI with various crystal targets during the last one and a half decade [33, 34, 6, 35, 36]. More recently, within the framework of the TECHNO-CLS project [13], a new 530 MeV positron beam line has been designed at MAMI [37] and the first experimental tests with the positron beam have been carried out [38, 39].
III.1 Statistical analysis of the channeling process
For each case study approximately $N_{0}=1.2\times 10^{4}$ of trajectories have been simulated and analyzed. Randomness in the ”entry conditions” leads to diversity in the scattering events for the different projectiles at the crystal entrance. Consequently, not all particles start their propagation moving in the the channeling regime. To quantify this property, the acceptance ${\cal A}$ can be introduced as the ratio of the number $N_{\rm acc}$ of particles accepted into the channeling regime to the total $N_{0}$ number of particles: ${\cal A}={N_{\rm acc}/N_{0}}$.
A significant methodological issue pertains to the formulation of a criterion for distinguishing between channeling and non-channeling regimes of projectiles’ motion. Depending on the theoretical approach employed to describe the interaction of a projectile with a crystalline environment, the criterion can be introduced in various ways. For instance, within the continuous potential framework, it is straightforward to define the channeling particles as those with transverse energies less than the height of the inter-planar (or inter-axial) potential barrier $\Delta U$. Within this framework, the acceptance is determined at the entrance to the crystal and can be defined as the ratio of the number of particles with transverse energy less than $\Delta U$ to the total number of particles. Within the framework of relativistic MD, the simulations are based on resolution of the EM (3), incorporating the interaction of a projectile particle with individual atoms of the crystal, as would be observed in reality. The potential experienced by the projectile varies rapidly during its motion, thereby coupling the transverse and longitudinal degrees of freedom. It is therefore necessary to propose an alternative criterion for identifying the channeling segments in the projectile’s trajectory. In the case of planar channeling, it can be postulated that a projectile is captured in the channeling mode when the sign of its transverse velocity undergoes a minimum of two changes within a single channel. [15].
The collisions (both elastic and inelastic ones) with the crystal atoms lead to a gradual decrease in the number of channeling particles, which were initially accepted, with the penetration distance. This dependence can be characterized by the fraction $f_{\rm ch,0}(z)=N_{\rm ch,0}(z)/N_{\rm acc}$ [19]. Here $N_{\rm ch,0}(z)$ is the number of particles which have been accepted at the entrance ($z=0$) and channeled to the distance $z$, where they dechannel.
The non-accepted particles, as well as the dechanneled particles, experience unrestricted over-barrier motion. However, due to the collisions, they can be captured in the channeling mode (rechanneled) at some point in the bulk. With the notation $N_{\rm ch}(z)$ as the total number of channeling particles at distance $z$, one can consider the fraction $f_{\rm ch}(z)=N_{\rm ch}(z)/N_{\rm acc}$, which accounts for the rechanneling process.
The fraction $f_{\rm ch,0}(z)$ can be used to quantify the dechanneling effect in terms of the penetration length $L_{\rm p}$, which stands for the mean length of the channeling segment of the trajectory of an accepted particle:
$L_{\rm p}=\int_{0}^{L}f_{\rm ch,0}(z){\rm d}z\,.$ (18)
As defined, $L_{\rm p}$ depends on the crystal length $L$. However, for sufficiently large values of $L$, for which $f_{\rm ch,0}(z)\to 0$ so that the contribution of the interval $[L,\infty]$ to the integral is negligibly small, Eq. (18) gives the value of $L_{\rm p}$ independent of $L$.
For both electrons and positrons, an analytical fit for $f_{\rm ch,0}(z)$ can be constructed using parameters calculated within the diffusion theory of the dechanneling process (see, e.g., [40, 2, 41]). Apart from some initial segment $[0,z_{0}]$, this fraction can be approximated by an exponentially decaying function:
For both electrons and positrons, an analytical fit can be constructed for $f_{\rm ch,0}(z)$ using parameters calculated within the diffusion theory of the dechanneling process (see e.g., [40, 2, 41]). Apart from some initial segment $[0,z_{0}]$, this fraction can be approximated by an exponentially decaying function:
$f_{\rm ch,0}(z)\approx\widetilde{f}_{\rm ch,0}(z)={\rm e}^{-{z-z_{0}\over\widetilde{L_{\rm d}}}}$ (19)
where $z_{0}$ and $L_{\rm d}$ are fitting parameters. It can be shown [19] that the quantities $L,$$L_{\rm p}$, $z_{0}$ and $\widetilde{L_{\rm d}}$ are related as follows: $L_{\rm p}=z_{0}+L_{\rm d}\left(1-{\rm e}^{-(L-z_{0})/L_{\rm d}}\right)$. In the case of a thick crystal, $L\gg L_{\rm d}$, one finds $L_{\rm p}\approx z_{0}+L_{\rm d}$.
Unlike $f_{\rm ch,0}(z)$, the dependence of the fraction $f_{\rm ch}(z)$ on the penetration distance is sensitive to the sign of the charge of the projectile. This will be discussed later in the paper.
Figures 1-3 and 4 show the dependence of the channeling fractions $f_{\rm ch,0}(z)$ (solid curves) and $f_{\rm ch}(z)$ (dashed curves) for electrons and positrons, respectively. The error bars indicate the statistical uncertainties due to the finite number of the simulated trajectories. In each graph, the fractions calculated with account for ionising collisions (marked ’yes’) are compared with those calculated without the ionising collisions.
Figure 1: Channeling fractions $f_{{\rm ch},0}$ (solid lines) and $f_{\rm ch}$ (dashed lines) versus penetration distance $z$ for 270 MeV electrons in diamond (left column) and silicon (right column) crystals. The upper row corresponds to the planar channeling along the (110) direction, the lower row - to the (111) direction. Curves with open circles show the results obtained taking into account the ionising collisions (labelled ’yes’ in the common legend placed in the upper right graph). Curves with filled circles represent the dependencies calculated without ionising collisions (labelled ’no’ in the legend). The simulations were performed for the beams of zero divergence.
Figure 2: Same as in Fig. 1 but for 855 MeV electrons.
Figure 3: Same as in Fig. 1 but (i) for 1500 MeV electrons, and (ii) adding the case of channeling in the (110) and (111) channels in germanium crystal, left column.
Let us first discuss the electron channeling, Figs. 1-3. The main feature seen in all graphs is that within the statistical errors the ’yes’ and ’no’ dependencies are virtually the same. The results presented indicate that for electrons the ionising collisions can be disregarded in the quantitative description of the dechanneling and rechanneling processes. This can be understood recalling that negatively charged particles experience channeling motion in the vicinity of the atomic planes. Therefore, they experience rather close collisions (i.e. with comparatively small impact parameter) with the crystal atoms. For these impact parameters, the change in the transverse momentum of the particle is mainly due to elastic scattering from the static atomic potential rather than inelastic scattering from the atomic electrons. The decrease in transverse momentum can cause the particle to be re-channeled and thus contribute to the channeling fraction $f_{\rm ch}(z)$.
As the penetration distance increases, the function $f_{\rm ch}(z)$ decreases much more slowly than the fraction $f_{\rm ch,0}(z)$. In Ref. [19] it was shown that $f_{\rm ch}(z)$ can be well approximated by the following fitting formula: